OpenStudy (anonymous):

. Solve . Write your solution using set builder notation. Please show all of your work. X+2>-3/11 im having trouble with this can anyone help?

OpenStudy (anonymous):

x+2>-3/11 or, x>-25/11

OpenStudy (anonymous):

\[x \in (-25/11,\infty)\]

OpenStudy (anonymous):

how do you write that in set builder notation?

OpenStudy (anonymous):

The range of x is from -25/11 (open) to infinity (open)

jimthompson5910 (jim_thompson5910):

Answer in set-builder notation is \[\Large \left\{x | x > -\frac{25}{11}\right\}\]

OpenStudy (anonymous):

thank you